OpenStudy (allnighter101):

f(x)= -3^2+2x-9.....Factor f(x) into a product of linear terms.

OpenStudy (mathstudent55):

Is your problem copied correctly? Should it be \(f(x) = -3x^2 + 2x - 9\) ? I don't like to work with a negative leading coefficient, so I'll start by factoring -1. \(f(x) = -3x^2 + 2x - 9\) \(f(x) = -1(3x^2 - 2x + 9)\) We need to factor a quadratic trinomial whose leading coefficient is not 1. The trinomial is of the form \(ax^2 + bx + c\). We can use the \(ac\) method. 1. Multiply \(ac\): \(3 \times 9 = 27\) 2. Factor \(ac\) into two factors whose sum is \(b\): there are no two such numbers, so this trinomial is not factorable.